# RingID: Rethinking Tree-Ring Watermarking for Enhanced Multi-Key Identification

### v2

@@ Appendix B Discarding Imaginary Part As Standalone Watermarking Approach
Relation with Tree-Ringrand [35] provides a variant called Tree-Ringrand that also injects noise as the watermark. However, they inject the same noise pattern for all generated images and intend to rely on pattern matching for watermark verification. The proposed method in this section distinguishes itself from Tree-Ringrand by injecting i.i.d. sampled noise for each generated image. The AUC for Tree-Ringrand and ours is 0.918 [35] and 0.901, respectively. The closely matched performances indicate that the deviation introduced by discarding imaginary part offers very robust discriminative power. This suggests that discarding imaginary part can effectively distinguish between watermarked and non-watermarked images even without relying on the specific noise pattern.
@@ Appendix C Failure Cases of Multi-Channel Rings
As discussed in the main text, we can imprint the ring watermarks onto multiple channels to trivially increase the capacity. However, we find that this often leads to the generation of ring-like artifacts, evident in a substantial proportion of cases, illustrated in Fig. 8. So we only imprint the ring watermark on a single channel by default.

### v3

@@ Appendix B Discarding Imaginary Part As Standalone Watermarking Approach
Relation with Tree-Ringrand [41] provides a variant called Tree-Ringrand that also injects noise as the watermark. However, they inject the same noise pattern for all generated images and intend to rely on pattern matching for watermark verification. The proposed method in this section distinguishes itself from Tree-Ringrand by injecting i.i.d. sampled noise for each generated image. The AUC for Tree-Ringrand [41] and ours is 0.918 and 0.901, respectively. The closely matched performances indicate that the deviation introduced by discarding imaginary part offers very robust discriminative power. This suggests that discarding imaginary part can effectively distinguish between watermarked and non-watermarked images even without relying on the specific noise pattern.
@@ Appendix C Failure Cases of Multi-Channel Rings
As discussed in the main text, we can imprint the ring watermarks onto multiple channels to increase the capacity. However, we find that this often leads to the generation of ring-like artifacts, evident in a substantial proportion of cases, illustrated in Fig. 8. So we only imprint the ring watermark on a single channel by default.
@@
Appendix D More Ablations and Comparisons
D.1 Results on Different Diffusion Models
D.2 Comparison with More Methods
Tab. 11 compares RingID with more methods on the watermark verification task. We can find that RingID achieves the best AUC and TPR@1%FPR under both clean and adversarial settings, showcasing strong robustness.
Table 11: Comparison with more methods on verification.
[…]
DwtDctSvd [9] 1.0 / 1.0 0.702 / 0.262 25.01 0.359
RivaGAN [49] 0.999 / 0.999 0.854 / 0.448 24.51 0.361
Tree-Ring [41] 1.0 / 1.0 0.975 / 0.694 25.93 0.364
RingID 1.0 / 1.0 0.995 / 0.926 26.13 0.365
Appendix E More Qualitatives
Fig. 9 gives more qualitative results.
Figure 9: More qualitative results. Images are generated by SD 2.1 with Stable-Diffusion-Prompts [18].